A ball (m = 0.1 kg) is thrown upwards with a velocity of 8 m/s. At what height will its velocity be 3 m/s?


1
Expert's answer
2022-06-10T08:28:47-0400
"h=\frac{v^2-v_0^2}{2a}\\
=\rm\frac{(3\: m/s)^2-(8\: m/s)^2}{2*(-9.8\: m/s^2)}=2.8\:\rm m"